# Polarization based direction of arrival estimation using a radio interferometric array

### v1

@@ 3 Transformer deep neural network
Further refinements such as residual connections and batch normalization $\mathrm{BN}(\cdot)$ (for stability while training) are appended,
${\bf X}^{\prime}_{\mathrm{out}}=\mathrm{BN}\left({\bf X}_{\mathrm{v}}+{\bf X}_{\mathrm{out}}\right)$ (18)
and furthermore, non-linear activation functions such as Gaussian error linear unit (GELU) activation are also used. Further such details are omitted for brevity but can be found in (Vaswani et al., 2017).
@@ 4 Simulations
The histogram of angular error for the DOA estimation using the trained DNN and using the minimum value of (14) for both arrays are shown in Figs. 7 and 9. Note that because of the inherent ambiguity, the angular error for SKA-Low array is calculated by considering both $\widehat{\phi}$ and $\widehat{\phi}+\pi$ as the possible estimates for the azimuth angle. Also noteworthy is the bias of the phase estimate (the peak of the histogram is not at zero), which is also reported in previous work Bell et al. (2025).
@@
We discuss the computational cost of the proposed method with respect to existing methods briefly. We consider the computational cost of the full eigenvalue decomposition of an $N\times N$ matrix to be $O(N^{3})$. We need to find the dominant eigenvector by performing the eigenvalue decomposition as in (9) where $N=4$, hence the complexity is $O(4^{3})$ (Note that we can use methods like the power iteration that have lower complexity, but we keep this larger complexity here). Therefore, the total complexity of running ESPRIT is $N_{l}\frac{N_{r}(N_{r}-1)}{2}O(4^{3})$ for the full set of sub-arrays. The other steps such as phase unwrapping and evaluating the DNN have fixed and lower computational complexity. In contrast, for running an algorithm like MUSIC (Schmidt, 1986), considering the full array to have $N_{l}N_{r}$ receivers, the complexity of the eigenvalue decomposition of the full correlation matrix is $O((N_{l}N_{r})^{3})$. Sparsity based methods (Yang et al., 2018; Zhang et al., 2024) need to create the dictionary matrix with $N_{g}^{2}$ or more steering vectors in 2D and will have matrices of size $(N_{l}N_{r})^{2}\times N_{g}^{2}$. Hence, compared to both these popular methods, the proposed method is much less computationally demanding.
@@ Acknowledgments
We dedicate this work to our passed colleague, Albert-Jan Boonstra.

### v2

@@ 2 The data model
Furthermore, it is also possible to formulate the cost function (14) to include the range $R$ as well as $f(R,\theta,\phi)$, and using this formulation to estimate the range in addition to the DOA for near-field sources is discussed in A.
@@ 3 Transformer deep neural network
For self-attention, when $N_{q}=N_{kv}$, further refinements such as residual connections and batch normalization $\mathrm{BN}(\cdot)$ (for stability while training) are appended. First we create a residual connection as
${\bf X}^{\prime}_{\mathrm{out}}=\mathrm{BN}\left({\bf X}_{\mathrm{v}}+{\bf X}_{\mathrm{out}}\right)\in\mathbb{R}^{N_{q}\times d_{\mathrm{model}}}$ (18)
and thereafter, non-linear activation is performed as
${\bf X}^{\prime\prime}_{\mathrm{out}}=\mathrm{BN}\left(\mathrm{GELU}\left({\bf X}^{\prime}_{\mathrm{out}}{\bf W}_{o1}\right){\bf W}_{o2}\right)\in\mathbb{R}^{N_{q}\times d_{\mathrm{model}}}$ (19)
where ${\bf W}_{o1}$ ($\in\mathbb{R}^{d_{\mathrm{model}}\times 4d_{\mathrm{model}}}$) and ${\bf W}_{o2}$ ($\in\mathbb{R}^{4d_{\mathrm{model}}\times d_{\mathrm{model}}}$) are additional learnable parameters. The Gaussian error linear unit activation is given by $\mathrm{GELU}(\cdot)$. The final output of the self-attention mechanism is ${\bf X}^{\prime\prime}_{\mathrm{out}}$.
@@ 4 Simulations
The histogram of angular error for the DOA estimation using the trained DNN and using the minimum value of (14) for both arrays are shown in Figs. 7 and 9. Note that because of the inherent ambiguity, the angular error for SKA-Low array is calculated by considering both $\widehat{\phi}$ and $\widehat{\phi}+\pi$ as the possible estimates for the azimuth angle. Also noteworthy is the bias of the phase estimate (the peak of the histogram is not at zero), which is also reported in previous work (Bell et al., 2025).
@@
We discuss the computational cost of the proposed method with respect to existing methods briefly. We consider the computational cost of the full eigenvalue decomposition of an $N\times N$ matrix to be $O(N^{3})$. We need to find the dominant eigenvector by performing the eigenvalue decomposition as in (9) where $N=4$, hence the complexity is $O(4^{3})$ (Note that we can use methods like the power iteration that have lower complexity, but we keep this larger complexity here). Therefore, the total complexity of running ESPRIT is $N_{l}\frac{N_{r}(N_{r}-1)}{2}O(4^{3})$ for the full set of sub-arrays. The other steps such as phase unwrapping and evaluating the DNN have fixed and lower computational complexity. In contrast, for running an algorithm like MUSIC (Schmidt, 1986), considering the full array to have $N_{l}N_{r}$ receivers, the complexity of the eigenvalue decomposition of the full correlation matrix is $O((N_{l}N_{r})^{3})$ and the search in $N_{g}^{2}$ grid points involve matrix-vector multiplication with complexity $O((N_{l}N_{r})^{2}N_{g}^{2})$. Sparsity based methods (Yang et al., 2018; Zhang et al., 2024) need to create the dictionary matrix with $N_{g}^{2}$ or more steering vectors in 2D and will have matrices of size $(N_{l}N_{r})^{2}\times N_{g}^{4}$ (because of the kronecker product). Hence, compared to both these popular methods, the proposed method is much less computationally demanding.
A quantitative comparison of the proposed method (AARTFAAC array) with MUSIC is presented in Fig. 11. We show the results for both far-field and near-field sources in this figure. For near-field sources, the MUSIC algorithm (using the full correlation matrix) gives better accuracy. The MUSIC algorithm is using steering vectors that are designed for a far-field source. Hence, as seen in Fig. 11 (b), when we have a source in the near-field of the array, the MUSIC algorithm designed for a far-field source completely fails to perform. In such a situation, both a direction and range search should be performed separately in MUSIC and the separation of the direction from the range is only possible for specialized array geometries and not for arbitrary arrays. In contrast, the proposed method gives accurate results both for far-field and near-field sources, without having to change the data model that is being used.
Figure 11: The comparison of the angular error of the transformer with MUSIC for 10000 samples. The histograms of the angular error for (a) the transformer DNN and (b) MUSIC (Schmidt, 1986) are shown. In each panel we show the angular error for sources in the far-field as well as in the near-field.
The computational time for the comparison shown in Fig. 11 was measured using a CPU with 8 threads. We measure the average time for one realization in the following results. The data used for the comparison were generated with 1 s duration and 1000 samples to form the correlation matrix. We consider the AARTFAAC array with $N_{r}=6$ and $N_{l}=48$ from Table 1. For the MUSIC algorithm, the time taken was 1.4 s (using 8 threads with the matrix size $2N_{l}N_{r}\times 2N_{l}N_{r}$). For the proposed method, the individual time measurements were as follows. For running ESPRIT on each sub-array (sequential, single thread, for $N_{r}(N_{r}-1)/2$ baselines) about 3e-4 s, consuming a total ($\times N_{l}$) time of 0.014 s. For phase unwrapping for each sub-array it consumed about 3e-4 s giving a total ($\times N_{l}$) time of 0.016 s. The evaluation of the DNN (transformer) took 0.004 s on the CPU. Therefore, the proposed algorithm took a total time of 0.034 s, which is much lower than the time taken by MUSIC.
@@ Acknowledgments
We dedicate this work to our passed colleague, Albert-Jan Boonstra. We also thank the anonymous reviewers for the careful review and valuable comments.
Appendix A Range and DOA estimation for near-field sources
For a near-field source, the source direction changes with respect to each baseline ${\bf b}_{pq}$. Therefore, we consider the centroid of each baseline to find the source direction vector $\widehat{\bf s}_{pq}$ as
$\widehat{\bf s}_{pq}=\frac{R\widehat{\bf s}-({\bf x}_{p}+{\bf x}_{q})/2}{\|R\widehat{\bf s}-({\bf x}_{p}+{\bf x}_{q})/2\|}$ (23)
and we formulate a cost function that also includes the range $R$ in addition to the DOA as
$f(R,\theta,\phi)=\sum_{pq}\left(\left(\widehat{\bf b}^{T}_{pq}\ \widehat{\bf s}_{pq}\right)^{2}-\left(\frac{\phi_{pq}}{2\pi b_{pq}/\lambda}\right)^{2}\right)^{2}.$ (24)
This function is evaluated on a grid of DOA $\overline{\theta}$$\overline{\phi}$ as well as on a discrete set of range values, $\overline{R}$ as $f(\overline{R},\overline{\theta},\overline{\phi})$. The transformer DNN in Fig. 2 is modified to not only output the DOA $\hat{\theta}$$\hat{\phi}$ but also the range $\log\hat{R}$. No additional input is fed to the DNN compared with the near-field case. In Fig. 12, we show an example where we train the transformer to not only estimate the DOA but also the range (AARTFAAC array). However, we only evaluate the range (log value) on 10 grid points uniformly sampled within the range given in Table 1, so this is a very coarse estimate. Nevertheless, we are able to get range results with reasonable accuracy albeit with a bias (range underestimated).
Figure 12: An illustration of the DOA and range estimation error of the transformer for the AARTFAAC array. The histograms of the (a) angular error and (b) range error are shown for 4000 examples of near-field sources.